Use the variational principle to obtain an upper limit to ground state energy of a particle in one dimensional box.


1
Expert's answer
2021-06-07T09:34:48-0400

"= <H>=<\psi/H/\psi>=\int_{-\infty}^{\infty}\psi*(H\psi)dz"

"H=\frac{eB_x}{m}Sy=\frac{e\bar{h}By}{m}[_1^0 \space _0^{-i}] /\psi>=[^{cos\theta}_{sin\theta}]"

So, "<\psi/=[cos\theta sin\theta]"

"=<H>=<\psi/H/\psi/>"

"=[cos \theta sin \theta][_1^0 \space_0^{-i}][_{sin\theta}^{cos \theta}] \frac{ehBy}{m}"

"=[-icos\theta sin\theta+sin \theta cos \theta]\frac{ehBy}{m}=0"

"\therefore <H>= <\psi/\mu/\psi>=0"